Fill in the blanks. Suppose the probability of a baseball player getting a hit in an at bat is 0.2935. If the player bats 24 times during a week, his number of hits should be around __________, give or take __________. Assume each at bat is independent.
Answer #1

Here, n = 24

p = 0.2935

q = 1 - p = 0.7065

Mean = np

= 24x0.2935

= 7.044

Standard deviation = \sqrt{npq}

= 2.2308

If the player bats 24 times during a week, his number of hits should be around 7.044, give or take 2.2308